OpenStudy (campayne):

Tell whether the product CD is defined for C9x5 and D5x9. If yes, give the product's dimensions as the following example: 2x3. If no, write no in the box

OpenStudy (unklerhaukus):

What are the inner dimensions of C_(9x5) D_(5x9) ?

OpenStudy (campayne):

ive no idea

OpenStudy (unklerhaukus):

\(\color{brown}{\text{these}}\) ones \[C_{9\times\color{brown}5}D_{\color{brown}5\times 9}\]

OpenStudy (campayne):

i dont know how to calculate the dimensions

OpenStudy (unklerhaukus):

the inner dimensions are both 5, so we can take the product of the matrices, the resulting matrix will have the outer dimensions

OpenStudy (campayne):

how do you do that

OpenStudy (unklerhaukus):

one way to think about it is that the 5's cancel and the 9's are left over. i.e. the resulting matrix will be 9x9

OpenStudy (campayne):

how do the 9s not cancel out each other

OpenStudy (unklerhaukus):

fortunately we do not have to calculate the element of this large (9 rows and 9 column ) matrix

OpenStudy (unklerhaukus):

the 9's don't cancel because they are on the outside

OpenStudy (campayne):

ohhhh i see

OpenStudy (unklerhaukus):

if we multiplied the matrices the other way around i.e. D_(5x9) C_(9x5) the result will be a 5 x 5 matrix

OpenStudy (campayne):

ahhh i see now

OpenStudy (campayne):

thanks c:

OpenStudy (unklerhaukus):

I that question before, with the two tables we were multiplying row and column vectors like this R_(1x3) P(3x1) = some (1x1) matrix i.e. the total

OpenStudy (campayne):

so its 1 x 1

OpenStudy (unklerhaukus):

yeah a number (a scalar) like 249.7 is a degenerate matrix

OpenStudy (campayne):

i see

OpenStudy (campayne):

wait a minute so how did you calculate all this